How do you solve 25x^4-61x^2+36=0?
Found 3 solutions by unlockmath, richwmiller, Alan3354:
Answer by unlockmath(1688)

Hello,
Let's pull out an x^2 like this: (25x^4-61x^2+36=0)
x^2(25x^2-61x+36)=0
This can be factored as:
x^2(25x-36)(x-1)=0 Solve x to get:
x=0
x=1
x=36/25
Plug in the answers to the first equation to see if it works.
Make Sense?
RJ
Check out a book I wrote:
www.math-unlock.com

Answer by richwmiller(17219)

This not finance. How lazy can you get?

Answer by Alan3354(69443)

How do you solve 25x^4-61x^2+36=0?
It's a quadratic in x^2
-----------------------
You can factor it:
(25x^2 - 36)*(x^2 - 1) = 0
x^2 = 1
x = +1, -1
-----------
25x^2 = 36
x = +6/5, -6/5